Calculate the pH of the resulting solution if 20.0 mL of 0.200 M HCl(aq) is added to 30.0 mL of 0.200 M NaOH(aq).

Calculate the pH of the resulting solution if 20.0 mL of 0.200 M HCl(aq) is added to 10.0 mL of 0.300 M NaOH(aq).
1
Expert's answer
2020-04-22T09:17:46-0400

(I) effective molarity is given by -

"MV=M_1V_1-M_2V_2"

Where


"M_1=0.2,M_2=0.2,V_1=30,V_2=20\ ml"

Now


"M(20+30)= 2\\M=0.04"

Molarity = "OH^-" ion concentration = 0.04


"pOH=log(\frac{1}{OH^-})=log(25)= 1.398"

pH= 14-pOH= 14-1.398=12.602


(II)

Similarly

"MV=M_1V_1-M_2V_2"


"M(20+10)=20\times0.2-10\times0.3\\M=0.033"

Here molarity = "H^+" concentration = 0.033


"pH= log(\frac{1}{H^+})= log(30)=1.477"